OpenStudy (anonymous):

find the inverse of f(x) = sqrt(4-e^2x)

OpenStudy (anonymous):

multiplicative inverse?

OpenStudy (anonymous):

1/sqrt(4-e^2x)

OpenStudy (anonymous):

thats not right

OpenStudy (anonymous):

solve it for x=... let f(x) = y then: y^2=4-e^2x 2x=ln|4-y^2| x=1/2 ln|4-y^2| now write inverse function as: f=1/2 ln|4-x^2|

OpenStudy (anonymous):

is it good?

OpenStudy (anonymous):

i dont understand the answer

OpenStudy (anonymous):

inverse function is 1/2 on natural ln of absolute value (...): \[f=1/2 \ln|4-x ^{2}|\]

OpenStudy (anonymous):

after y^2=4-e^2x I am LOST!

OpenStudy (anonymous):

how do you finish solving for x

OpenStudy (anonymous):

e^2x=4-y^2 take ln both sides: 2x=ln|4-y^2| x=... because it's inverse - just change places for x& y

OpenStudy (anonymous):

ok i understand now

OpenStudy (anonymous):

so what will be the domain of the inverse function

OpenStudy (anonymous):

\[4-x^{2}>0\] because ln

OpenStudy (anonymous):

\[x^{2}<4\] -2<x<2

OpenStudy (anonymous):

-2 doesnt work

OpenStudy (anonymous):

that's why (-2) & 2 are not included. it simple more & less (not equal)